Question 1: Talal Industries developed the following information for the product itsells: of total Sales Variable cost of goo please provide step by step answer prefairable handwritten
Answer #1

a)

Sales(100000*50) 5000000
Less:
Variable COGS (23*100000) 2300000
Selling costs (5000000*10%) 500000
Administrative costs (100000*2) 200000
Contribution margin 2000000
Less:
Fixed COGS 800000
Fixed selling 500000
Fixed administrative 300000
Profit 400000

b)

Contribution margin per unit = 2000000/100000 = 20

Contribution margin ratio = (20/50)*100 = 40%

c)

Breakeven point(units) = Fixed cost/Contribution per unit = (800000+500000+300000)/20 = 80000 units

Breakeven point ($) = 80000 units * 50 = 4000000

d)

Margin of safety = (100000-80000)/100000 = 20%

Degree of operating leverage = Contribution margin / (Contribution margin-Fixed costs)

= (5000000-2300000-500000-200000) / (5000000-2300000-500000-200000-800000-500000-300000)

= 2000000/400000

= 5 times